OpenStudy (anonymous):

The sum of a quantity "p" together with its two-third, and its half becomes 234. What is the quantity "p"?

OpenStudy (anonymous):

\[p+\frac{2}{3}p+\frac{1}{2}p=334\]

OpenStudy (anonymous):

\[p+\frac{2}{3}p+\frac{1}{2}p=234\]

OpenStudy (anonymous):

combine like terms on the left, get \[\frac{13}{6}p=234\] then \[p=234\times \frac{6}{13}\] whatever that is

OpenStudy (anonymous):

thank you i got p=108